# Magnetic disk winds in protoplanetary disks:

### v2

@@ Abstract
Context. Canonically, a protoplanetary disk is thought to undergo (gravito-)viscous evolution, wherein the angular momentum of the accreting material is transported outwards. However, several lines of reasoning suggest that the turbulent viscosity in a typical protoplanetary disk is insufficient to drive the observed accretion rates. An emerging paradigm suggests that radially extended magnetic disk winds, which transport angular momentum vertically, may play a crucial role in the disk evolution.
Aims. We propose a global model of magnetic wind-driven accretion for evolution of protoplanetary disks in the thin-disk limit, based on the insights gained from local shearing box simulations. In this paper, we aim to develop this model and constrain the model parameters with the help of theoretical expectations and comparison with observations.
Methods. The magnetic wind is characterized with the associated loss of angular momentum and mass, which are modeled with fitting formulae that depend on the local disk conditions and stellar properties. We incorporate the disk winds self-consistently in the numerical magnetohydrodynamic code FEOSAD and study formation and long-term evolution of protoplanetary disks. We include disk self-gravity and an adaptive turbulent $\alpha$ that depends on the local ionization balance, while the co-evolution of a two part dusty component is also considered. Synthetic observations are obtained via detailed modeling with radiation thermo-chemical code ProDiMo.
Results. The models with inclusion of disk winds satisfy general expectations from both theory and observations. The disk wind parameters can be guided by observational constraints, and the synthetic observations resulting from such a model compare favorably with the selected ALMA survey data of Class II disks. The proposed magnetic disk wind model is a significant step forward in the direction of representing a more complete disk evolution, wherein the disk experiences concurrent torques from viscous, gravitational, and magnetic wind processes.
@@ 1 Introduction
During its formation, a low mass star accretes through a flattened, centrifugally-supported circumstellar disk. Such a disk inevitably forms due to the angular momentum contained in the parent molecular cloud or through local subsonic turbulence (Yorke et al., 1993; Li et al., 2014; Verliat et al., 2020). For accretion of gas to proceed through the disk, the accreting material must efficiently lose its angular momentum. Gravitational instability (GI) facilitates angular momentum redistribution during protostellar disk formation from a magnetized prestellar core (Tsukamoto et al., 2015; Masson et al., 2016; Xu & Kunz, 2021; Mauxion et al., 2024), and continues to play a dominant role during the early embedded phases of the disk evolution (Vorobyov & Basu, 2006, 2009; Kratter & Lodato, 2016; Rice, 2016). This occurs when the disk is sufficiently massive as compared to the host star, and the self-gravity of the disk leads to formation of spiral arms, density fluctuations and fragmentation, resulting in efficient global mass transport.
@@
Later in the Class II stage, a protoplanetary disk (PPD) is canonically thought to evolve viscously, with magnetorotational instability (MRI) providing the turbulence necessary for this viscosity (Shakura & Sunyaev, 1973; Pringle, 1981; Balbus & Hawley, 1991; Turner et al., 2014). When a weak magnetic field embedded in the disk interacts with the partially ionized material, the Keplerian shear in the disk gives rise to MRI. The action of MRI requires sufficiently ionized gas, and cosmic rays are thought to be the primary source of ionization in PPDs (Umebayashi & Nakano, 1981, 1988; Thi et al., 2019). The cosmic rays penetrate the disk surface from the outside, passing through the entire disk thickness at large radii. The innermost regions have high enough surface density that the cosmic rays cannot reach the midplane, resulting in insufficient ionization to couple to the magnetic field. This suppresses the MRI and causes the PPD to accrete through a layered structure, wherein the accretion occurs only through the ionized surface layers and a magnetically dead zone is formed at the midplane (Gammie, 1996). The dead zone forms a bottleneck for mass and angular momentum transport and gives rise to disk substructure such as rings and vortices, which form pressure maxima and sites for dust growth (Dzyurkevich et al., 2010; Pinilla et al., 2012; Kadam et al., 2022; Regály et al., 2023). Spatially resolved images of PPDs indeed show an abundance of disk substructure in the majority of large, bright disks (Andrews et al., 2009; Flock et al., 2015; Andrews et al., 2018; Zhang et al., 2018; Law et al., 2021; Hsieh et al., 2024).
@@
However, several lines of reasoning point towards a dramatically different picture of PPD evolution, at least in the Class II stage. The Shakura-Sunyaev $\alpha$-parameter quantifies the efficiency of angular momentum transport in the standard theory of viscous accretion (Shakura & Sunyaev, 1973). In the absence of a dead zone, accretion within the disk is most efficient and a typical value of $\alpha$ is about 0.01 (Hartmann et al., 1998; Hughes et al., 2011; Andrews et al., 2010). However, recent observations of protoplanetary disks indicate that the inferred value of the $\alpha$ parameter is much lower, which is insufficient for the disk to accrete within its typical observed lifespan of a few Myr (Williams & Cieza, 2011). For example, the dust rings observed within PPDs by Atacama Large Millimeter/ Submillimeter Array (ALMA) are significantly thin, and studies of dust properties suggest the turbulence to be $\alpha\sim 10^{-3}-10^{-4}$ (Pinte et al., 2016; Dullemond et al., 2018; Muro-Arena et al., 2018; Franceschi et al., 2023). However, some uncertainty still exists from the assumption of dust properties, since the dusty rings can only constrain the ratio of $\alpha$ to the Stokes number. Spatially and spectrally resolved molecular line measurements also indicate that the contribution of turbulence to the line broadening is consistent with weak turbulence (Flaherty et al., 2015, 2017, 2020). Conversely, from a theoretical viewpoint, the disks in recent magnetohydrodynamic (MHD) simulations also point towards marginal excitation of MRI turbulence. When the non-ideal MHD effects, especially ambipolar diffusion, are included, the simulations show increasingly laminar flow and much less overall turbulence. Instead, magnetic disk winds are launched near the disk surface, which carry a significant amount of angular momentum (Suzuki & Inutsuka, 2009; Bai & Stone, 2013; Fromang et al., 2013; Simon et al., 2013; Béthune et al., 2017; Zhu & Stone, 2018; Cui & Bai, 2020; Lesur, 2021b).
@@
The mechanism for angular momentum loss via disk winds was first described by Blandford & Payne (1982a), and it is similar to magnetic braking of rapidly rotating stars. When the disk magnetic field makes a sufficiently large angle with the normal ($>30^{\circ}$), the outflowing gas, which is constrained to move along the open magnetic field lines, is accelerated centrifugally. The gas flow eventually bends the poloidal magnetic field component into a toroidal field, and the gradient of the toroidal magnetic pressure assists in the collimation of the wind (Pudritz & Ray, 2019). We note that these magnetocentrifugal winds are fundamentally different from photoevaporative winds; the latter are launched thermally from the disk surface due to external irradiation and do not exert a torque on the disk (Hollenbach et al., 1994; Alexander et al., 2014; Ercolano & Pascucci, 2017). Photoevaporation is negligible in the early disk evolution, although it becomes significant in the late stages, when the stellar accretion rate becomes comparable to photoevaporation rate. Photoevaporative winds are considered to be responsible for the eventual dispersal of the gas in PPDs (Alexander et al., 2006, 2014). In reality, these two mass ejection processes operate simultaneously and can be collectively considered as a “magneto-thermal wind” (Bai et al., 2016; Lesur et al., 2023).
@@
High spectral resolution observations of gas lines that trace unbound disk material (forbidden lines such as [OI] $\lambda 6300$ and [SII] $\lambda 6731$) show the ubiquitous presence of outflows around PPDs (Pascucci et al., 2023). These outflows are divided into a high-velocity component (HVC), traveling at hundreds of ${\rm km\,s^{-1}}$ and a low velocity component (LVC) of 10-50 ${\rm km\,s^{-1}}$ (Natta et al., 2014; McGinnis et al., 2018; Nisini et al., 2018; Rab et al., 2022; Fang et al., 2023). The HVC is interpreted as a collimated axial jet, which is launched very close to the surface of the accreting object via magnetic processes. The LVC component originating beyond the dust evaporation radius and spanning a majority of the disk is typically identified as magnetocentrifugal wind, although it may also be interpreted as X-ray-driven photoevaporative flow (Rab et al., 2023). Spatially resolved outflows of young objects show small-scale rotating flows as slow molecular winds, in addition to fast moving jets (Güdel et al., 2018; de Valon et al., 2020; Lee et al., 2021). These winds show a nested onion-like morphology at large opening semi-angles (20-45) and may originate within the inner 40 au of the disk. The mass flux estimates of the disk winds are exceedingly significant, with 1-10 times the jet mass flux and 0.3-3 times the accretion rate onto the central protostar (Tabone et al., 2020). Thus, an emerging paradigm of protoplanetary disk evolution suggests that in addition to redistribution of angular momentum through gravitoviscous torques, extraction of mass and angular momentum via magnetic disk winds plays a crucial role as a driver of accretion.
@@
As the winds fundamentally arise from the interaction of partially ionized gas with the magnetic field, theoretical insights into disk winds primarily come from 3D non-ideal MHD simulations. This includes both local shearing box studies (Bai & Stone, 2013; Simon et al., 2013; Suzuki & Inutsuka, 2009) and global simulations which resolve a significant fraction of the disk (Béthune et al., 2017; Zhu & Stone, 2018; Cui & Bai, 2020; Lesur, 2021b). Since long integration times of the order of the disk lifetime (few Myr) is not feasible with 3D simulations, disk winds can be studied with simplified semi-analytical global evolutionary models in 1D (Kunitomo et al., 2020; Alessi & Pudritz, 2022; Somigliana et al., 2022; Tabone et al., 2022). However, certain important processes occurring in PPDs cannot be faithfully represented in 1D axisymmetry. For example, in the early stages of disk evolution, GI leads to formation of large-scale spirals and clumps, which may eventually form wide orbit gas giants (Boss, 1997; Vorobyov & Basu, 2010a; Kratter & Lodato, 2016). In the later stages of evolution, disk-planet interactions are also intrinsically axially asymmetric (Paardekooper et al., 2023). Thus, theoretical studies of long-term accretion disk evolution are often performed in the plane of the disk ($r,\phi$) in 2D, with the thin-disk approximation. The vertical hydrostatic equilibrium is imposed throughout the disk, so that the vertical temperature gradients and motions are neglected. Such an approximation typically works very well for PPDs, since the ratio of gas scale height to the local radius ($H_{\rm g}/r$) is much less than unity. This approach is powerful for studying long-term behavior of PPDs, because it relaxes many of the simplifying assumptions of 1D Pringle (1981)-like viscous equation of disk evolution, while also being computationally feasible.
@@
Almost all of the thin-disk models of protoplanetary disks to date assume a viscous evolution of the disk; the disk self gravity is often considered, however, magnetic winds are essentially neglected. Recently, Kimmig et al. (2020) and Elbakyan et al. (2022) have considered the effects of wind-driven accretion in the thin-disk limit, in the context of planetary migration and gap opening, respectively. Nevertheless, these models leave much room for improvement. The wind parameters such as its escape velocity and the angle of the magnetic field cannot be held constant over the disk lifetime. Similarly, the vertical angular momentum transport of magnetic disk winds is fundamentally different from the Shakura & Sunyaev (1973) $\alpha$ parameter and cannot be modeled as an additional viscosity, as the latter only redistributes the angular momentum within the disk. In addition, the co-evolution of the dust component is not explicitly considered. Furthermore, accretion variability during star formation complicates the situation, especially the powerful FUor-type outbursts, wherein the accretion luminosity of the protostar increases by hundreds of times (Hartmann & Kenyon, 1996; Audard et al., 2014; Fischer et al., 2023). These events are important over the long term because they can significantly alter the disk composition, dynamics, and snow lines, while as much as 10-35% of the main-sequence stellar mass can be accreted via episodic accretion (Offner & McKee, 2011; Dunham & Vorobyov, 2012; Fischer et al., 2019; Molyarova et al., 2018; Vorobyov et al., 2020a, 2022). The accretion variability during these events should also directly lead to a corresponding variability in the wind outflow rate (Konigl & Pudritz, 2000; Ellerbroek et al., 2014). With these issues in mind, a model framework of magnetic disk winds for protoplanetary disks in the thin-disk limit, which incorporates all of the aforementioned phenomena, is desirable.
@@
Here, we propose a new global model of magnetic disk winds that takes into account the insights gained from local, shearing box MHD simulations. The results of shearing box simulations conducted at several radial distances can be stitched together in order to estimate the wind behavior throughout the PPD (Bai, 2013; Bai & Stone, 2013). We improve upon the wind mass loss rate and surface stress estimates of Bai et. al. by taking into account the corrections that are essential for long-term disk evolution. The resulting prescription of winds is incorporated self-consistently into the global MHD code FEOSAD, which is capable of simulating the formation and evolution of PPDs over long timescales in the thin-disk limit (Vorobyov et al., 2020b). Thus, all three major sources of angular momentum transport are now included; namely, turbulent viscosity, gravitational instability, and magnetic disk winds. We follow the disk evolution from the earliest embedded Class 0/I stages through the later Class II stage, demonstrating that the proposed model of magnetic disk winds is a reasonable approximation in the thin-disk limit. We confirm that the disk winds have significant impact on the disk evolution and global disk properties, most notably, disks with winds tend to be smaller in both size and mass. The simulation results are post-processed with the radiation thermo-chemical disk code ProDiMo (Woitke et al., 2009, 2016) and the obtained synthetic observations are compared with ALMA survey data of Class II disks. Our analysis demonstrates that such a comparison can be used to guide and constrain the disk wind model parameters.
@@
In the interest of conciseness and readability of the study, we present in this paper our disk wind model in detail, and explain the steps of obtaining synthetic observations and comparison with large-scale surveys. In a second companion paper, we will comprehensively investigate the properties of magnetic disk winds, e.g., characteristic bulk and radial properties of the winds, their effects on substructure formation, and comparison of different modes of angular momentum transport. The structure of the current paper is as follows. Section 2.1 describes our MHD disk evolution code and Section 2.2 elaborates on the magnetic wind model. In Section 2.3 we explain how the synthetic observations are obtained, while Section 2.4 describes the initial conditions. In Section 3, the results of the numerical simulations as well as radiation thermo-chemical modeling are presented in the context of observational constraints and theoretical consensus. The findings are summarized in Section 4.

### v3

@@ Abstract
Context. Canonically, a protoplanetary disk is thought to undergo (gravito-)viscous evolution wherein the angular momentum of the accreting material is transported outward. However, several lines of reasoning suggest that the turbulent viscosity in a typical protoplanetary disk is insufficient to drive the observed accretion rates. An emerging paradigm suggests that radially extended magnetic disk winds, which transport angular momentum vertically, may play a crucial role in disk evolution.
Aims. We propose a global model of magnetic wind-driven accretion for the evolution of protoplanetary disks in the thin-disk limit based on the insights gained from local shearing box simulations. In this paper, we aim to develop this model and constrain the model parameters with the help of theoretical expectations and through comparison with observations.
Methods. The magnetic wind is characterized with the associated loss of angular momentum and mass, and we modeled these with fitting formulae that depend on the local disk conditions and stellar properties. We incorporated the disk winds self-consistently in the numerical magnetohydrodynamic code FEOSAD and studied the formation and long-term evolution of protoplanetary disks. We included disk self-gravity and an adaptive turbulent $\alpha$ that depends on the local ionization balance, while the co-evolution of a two-part dusty component was also considered. We obtained synthetic observations via detailed modeling with the radiation thermo-chemical code ProDiMo.
Results. The models that include disk winds satisfy the general expectations from both theory and observations. The disk wind parameters can be guided by observational constraints, and the synthetic observations resulting from such a model compare favorably with the selected ALMA survey data of Class II disks. The proposed magnetic disk wind model is a significant step forward in the direction of representing a more complete disk evolution, wherein the disk experiences concurrent torques from viscous, gravitational, and magnetic wind processes.
@@ 1 Introduction
During its formation, a low mass star accretes through a flattened centrifugally supported circumstellar disk. Such a disk inevitably forms due to the angular momentum contained in the parent molecular cloud or through local subsonic turbulence (Yorke et al., 1993; Li et al., 2014; Verliat et al., 2020). For the accretion of gas to proceed through the disk, the accreting material must efficiently lose its angular momentum. Gravitational instability (GI) facilitates angular momentum redistribution during protostellar disk formation from a magnetized prestellar core (Tsukamoto et al., 2015; Masson et al., 2016; Xu & Kunz, 2021; Mauxion et al., 2024) and continues to play a dominant role during the early embedded phases of the disk’s evolution (Vorobyov & Basu, 2006, 2009; Kratter & Lodato, 2016; Rice, 2016). This occurs when the disk is sufficiently massive as compared to the host star, and the self-gravity of the disk leads to the formation of spiral arms, density fluctuations, and fragmentation, resulting in efficient global mass transport.
@@
Later in the Class II stage, a protoplanetary disk (PPD) is canonically thought to evolve viscously, with magnetorotational instability (MRI) providing the turbulence necessary for this viscosity (Shakura & Sunyaev, 1973; Pringle, 1981; Balbus & Hawley, 1991; Turner et al., 2014). When a weak magnetic field embedded in the disk interacts with the partially ionized material, the Keplerian shear in the disk gives rise to MRI. The action of MRI requires sufficiently ionized gas, and cosmic rays are thought to be the primary source of ionization in PPDs (Umebayashi & Nakano, 1981, 1988; Thi et al., 2019). The cosmic rays penetrate the disk surface from the outside, passing through the entire disk thickness at large radii. The innermost regions have a high enough surface density that the cosmic rays cannot reach the midplane, resulting in insufficient ionization to couple to the magnetic field. This suppresses the MRI and causes the PPD to accrete through a layered structure, wherein the accretion occurs only through the ionized surface layers, and a magnetically dead zone is formed at the midplane (Gammie, 1996). The dead zone forms a bottleneck for mass and angular momentum transport and gives rise to such disk substructures as rings and vortices, which form pressure maxima and sites for dust growth (Dzyurkevich et al., 2010; Pinilla et al., 2012; Kadam et al., 2022; Regály et al., 2023). Spatially resolved images of PPDs indeed show an abundance of disk substructure in the majority of large, bright disks (Andrews et al., 2009; Flock et al., 2015; Andrews et al., 2018; Zhang et al., 2018; Law et al., 2021; Hsieh et al., 2024).
@@
However, several lines of reasoning point toward a dramatically different picture of PPD evolution, at least in the Class II stage. The Shakura-Sunyaev $\alpha$-parameter quantifies the efficiency of angular momentum transport in the standard theory of viscous accretion (Shakura & Sunyaev, 1973). In the absence of a dead zone, accretion within the disk is most efficient, and a typical value of $\alpha$ is about 0.01 (Hartmann et al., 1998; Hughes et al., 2011; Andrews et al., 2010). However, recent observations of protoplanetary disks indicate that the inferred value of the $\alpha$ parameter is much lower, which is insufficient for the disk to accrete within its typical observed lifespan of a few million years (Williams & Cieza, 2011). For example, the dust rings observed within PPDs by the Atacama Large Millimeter/ Submillimeter Array (ALMA) are significantly thin, and studies of dust properties suggest the turbulence to be $\alpha\sim 10^{-3}-10^{-4}$ (Pinte et al., 2016; Dullemond et al., 2018; Muro-Arena et al., 2018; Franceschi et al., 2023). However, some uncertainty still exists from the assumption of dust properties since the dusty rings can only constrain the ratio of $\alpha$ to the Stokes number. Spatially and spectrally resolved molecular line measurements also indicate that the contribution of turbulence to the line broadening is consistent with weak turbulence (Flaherty et al., 2015, 2017, 2020). Conversely, from a theoretical viewpoint, the disks in recent magnetohydrodynamic (MHD) simulations also point toward marginal excitation of MRI turbulence. When the non-ideal MHD effects are included, especially ambipolar diffusion, the simulations show an increasingly laminar flow and much less overall turbulence. Instead, magnetic disk winds are launched near the disk surface, which carry a significant amount of angular momentum (Suzuki & Inutsuka, 2009; Bai & Stone, 2013; Fromang et al., 2013; Simon et al., 2013; Béthune et al., 2017; Zhu & Stone, 2018; Cui & Bai, 2020; Lesur, 2021b).
@@
The mechanism for angular momentum loss via disk winds was first described by Blandford & Payne (1982a), and it is similar to magnetic braking of rapidly rotating stars. When the disk magnetic field makes a sufficiently large angle with the normal ($>30^{\circ}$), the outflowing gas, which is constrained to move along the open magnetic field lines, is accelerated centrifugally. The gas flow eventually bends the poloidal magnetic field component into a toroidal field, and the gradient of the toroidal magnetic pressure assists in the collimation of the wind (Pudritz & Ray, 2019). We note that these magnetocentrifugal winds are fundamentally different from photoevaporative winds; the latter are launched thermally from the disk surface due to external irradiation and do not exert a torque on the disk (Hollenbach et al., 1994; Alexander et al., 2014; Ercolano & Pascucci, 2017). Photoevaporation is negligible in the early disk evolution, although it becomes significant in the late stages, when the stellar accretion rate becomes comparable to the photoevaporation rate. Photoevaporative winds are considered to be responsible for the eventual dispersal of the gas in PPDs (Alexander et al., 2006, 2014). In reality, these two mass ejection processes operate simultaneously and can be collectively considered as a “magneto-thermal wind” (Bai et al., 2016; Lesur et al., 2023).
@@
High spectral resolution observations of gas lines that trace unbound disk material (forbidden lines such as [OI] $\lambda 6300$ and [SII] $\lambda 6731$) show the ubiquitous presence of outflows around PPDs (Pascucci et al., 2023). These outflows are divided into a high-velocity component (HVC) traveling at hundreds of kilometers per second and a low velocity component (LVC) of 10-50 ${\rm km\,s^{-1}}$ (Natta et al., 2014; McGinnis et al., 2018; Nisini et al., 2018; Rab et al., 2022; Fang et al., 2023). The HVC is interpreted as a collimated axial jet, which is launched very close to the surface of the accreting object via magnetic processes. The LVC component originating beyond the dust evaporation radius and spanning a majority of the disk is typically identified as magnetocentrifugal wind, although it may also be interpreted as X-ray driven photoevaporative flow (Rab et al., 2023). Spatially resolved outflows of young objects show small-scale rotating flows as slow molecular winds in addition to fast moving jets (Güdel et al., 2018; Lee et al., 2021). These winds show a nested onion-like morphology at large opening semi-angles (20-50) and may originate within the inner ten au of the disk (de Valon et al., 2020; Pascucci et al., 2025). The mass flux estimates of the disk winds are exceedingly significant, with one to ten times the jet mass flux and 0.3 to three times the accretion rate onto the central protostar (Tabone et al., 2020). Thus, an emerging paradigm of protoplanetary disk evolution suggests that in addition to redistribution of angular momentum through gravitoviscous torques, extraction of mass and angular momentum via magnetic disk winds plays a crucial role as a driver of accretion.
@@
As the winds fundamentally arise from the interaction of partially ionized gas with the magnetic field, theoretical insights into disk winds primarily come from 3D non-ideal MHD simulations. This includes both local shearing box studies (Bai & Stone, 2013; Simon et al., 2013; Suzuki & Inutsuka, 2009) and global simulations, which resolve a significant fraction of the disk (Béthune et al., 2017; Zhu & Stone, 2018; Cui & Bai, 2020; Lesur, 2021b). Since long integration times of the order of the disk lifetime (a few million years) are not feasible with 3D simulations, disk winds can be studied with simplified semi-analytical global evolutionary models in 1D (Kunitomo et al., 2020; Alessi & Pudritz, 2022; Somigliana et al., 2022; Tabone et al., 2022). However, certain important processes occurring in PPDs cannot be faithfully represented in 1D axisymmetry. For example, in the early stages of disk evolution, GI leads to the formation of large-scale spirals and clumps, which may eventually form wide orbit gas giants (Boss, 1997; Vorobyov & Basu, 2010a; Kratter & Lodato, 2016). In the later stages of evolution, disk-planet interactions are also intrinsically axially asymmetric (Paardekooper et al., 2023). Thus, theoretical studies of long-term accretion disk evolution are often performed in the plane of the disk ($r,\phi$) in 2D, with the thin-disk approximation. The vertical hydrostatic equilibrium is imposed throughout the disk so that the vertical temperature gradients and motions are neglected. Such an approximation typically works very well for PPDs since the ratio of gas scale height to the local radius ($H_{\rm g}/r$) is much less than unity. This approach is powerful for studying the long-term behavior of PPDs because it relaxes many of the simplifying assumptions of the 1D Pringle (1981)-like viscous equation of disk evolution while also being computationally feasible.
@@
Almost all of the thin-disk models of protoplanetary disks to date assume a viscous evolution of the disk. The disk self-gravity is often considered, but magnetic winds are essentially neglected. Recently, Kimmig et al. (2020) and Elbakyan et al. (2022) have considered the effects of wind-driven accretion in the thin-disk limit in the context of planetary migration and gap opening, respectively. Nevertheless, these models leave much room for improvement. The wind parameters, such as its escape velocity and the angle of the magnetic field, cannot be held constant over the disk lifetime. Similarly, the vertical angular momentum transport of magnetic disk winds is fundamentally different from the Shakura & Sunyaev (1973) $\alpha$ parameter and cannot be modeled as an additional viscosity, as the latter only redistributes the angular momentum within the disk. In addition, the co-evolution of the dust component is not explicitly considered. Furthermore, accretion variability during star formation complicates the situation, especially the powerful FUor-type outbursts, wherein the accretion luminosity of the protostar increases by hundreds of times (Hartmann & Kenyon, 1996; Audard et al., 2014; Fischer et al., 2023). These events are important over the long term because they can significantly alter the disk composition, dynamics, and snow lines, while as much as 10-35% of the main-sequence stellar mass can be accreted via episodic accretion (Offner & McKee, 2011; Dunham & Vorobyov, 2012; Fischer et al., 2019; Molyarova et al., 2018; Vorobyov et al., 2020a, 2022). The accretion variability during these events should also directly lead to a corresponding variability in the wind outflow rate (Konigl & Pudritz, 2000; Ellerbroek et al., 2014). With these issues in mind, a model framework of magnetic disk winds for protoplanetary disks in the thin-disk limit, which incorporates all of the aforementioned phenomena, is desirable.
@@
In this work, we propose a new global model of magnetic disk winds that takes into account the insights gained from local shearing box MHD simulations. The results of the shearing box simulations conducted at several radial distances can be stitched together in order to estimate the wind behavior throughout the PPD (Bai, 2013; Bai & Stone, 2013). We improve upon the wind mass loss rate and surface stress estimates of Bai et. al. by taking into account the corrections that are essential for long-term disk evolution. The resulting prescription of winds is incorporated self-consistently into the global MHD code FEOSAD, which is capable of simulating the formation and evolution of PPDs over long timescales in the thin-disk limit (Vorobyov et al., 2020b). Thus, all three major sources of angular momentum transport are now included, namely, turbulent viscosity, gravitational instability, and magnetic disk winds. We follow the disk evolution from the earliest embedded Class 0 or I stages through to the later Class II stage, demonstrating that the proposed model of magnetic disk winds is a reasonable approximation in the thin-disk limit. We confirm that the disk winds have a significant impact on the disk evolution and global disk properties. Most notably, the disks with winds tend to be smaller in both size and mass. The simulation results are post-processed with the radiation thermo-chemical disk code ProDiMo (Woitke et al., 2009, 2016), and the obtained synthetic observations are compared with ALMA survey data of Class II disks. Our analysis demonstrates that such a comparison can be used to guide and constrain the disk wind model parameters.
@@
In the interest of the conciseness and readability of the study, we present in this paper our disk wind model in detail and explain the steps of obtaining synthetic observations and the subsequent comparison with large-scale surveys. In a second companion paper, we will comprehensively investigate the properties of magnetic disk winds, for example, characteristic bulk and radial properties of the winds, their effects on substructure formation, and comparison of different modes of angular momentum transport. The structure of the current paper is as follows. Section 2.1 describes our MHD disk evolution code, and Sect. 2.2 elaborates on the magnetic wind model. In Sect. 2.3, we explain how the synthetic observations are obtained, while Sect. 2.4 describes the initial conditions. In Sect. 3, the results of the numerical simulations as well as radiation thermo-chemical modeling are presented in the context of observational constraints and theoretical consensus. The findings are summarized in Sect. 4.